OpenStudy (anonymous):

How do I compute the double integral over S of x+y (posted in equation editor) where S is the region in the first quadrant (x>0, y>0) inside of the disk x^2 + y^2 <= a^2 and under the line y=sqrt(3)x.

OpenStudy (anonymous):

I figure that one should convert to polar form integrating: \[\int\limits_{0}^{\pi/2}\int\limits_{0}^{a/2}r^2(\cos\theta + \sin\theta) dr d\theta\] I got a/2 from substituting y=sqrt(3)x into x^2+y^2=a^2

OpenStudy (anonymous):

what is sqrt(3)x supposed to be? sqrt(3x)? cube root of 3x?

OpenStudy (anonymous):

cube root of 3 multiplied by x

OpenStudy (anonymous):

cube root of 3 times x? (3)^(1/3) * x?

OpenStudy (anonymous):

sorry square root of 3 so: \[\sqrt{3} \times x\]

OpenStudy (anonymous):

well theta isn't from 0 to pi/2

OpenStudy (anonymous):

it should be from 0 to pi/3. and radius is from 0 to a

OpenStudy (anonymous):

Right, I understand now that theta isn't 0 to pi/2 but how did you arrive at pi/3?

OpenStudy (anonymous):

tan(theta) = y/x = sqrt(3)x/x = sqrt(3) tan(pi/3) = sqrt(3)

OpenStudy (anonymous):

Thanks, didn't think of that!